40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 26 01:00:23 2017, gautam, Update, General, DRMI locking  DRMI_1f_26April2017.pngIMG_7373.JPGIMG_7372.JPG
    Reply  Fri Apr 28 02:04:36 2017, gautam, Update, General, DRMI locking  20170428_DRMI.pdf
       Reply  Fri Apr 28 13:56:26 2017, rana, Update, General, DRMI locking  
          Reply  Mon May 1 16:29:51 2017, gautam, Update, General, DRMI locking  LSC_sensingNoise.pdfIMG_7378.JPG
             Reply  Mon May 1 21:45:54 2017, ericq, Update, General, DRMI locking  
                Reply  Wed May 3 18:45:45 2017, rana, Update, General, DRMI locking  
                   Reply  Thu May 4 19:03:15 2017, gautam, Update, General, DRMI locking - preliminary MICH NB DRMI_noArms_April30.pdfMICH_OLTF.pdfC1NB_disp_40m_MICH_NB_30_April_2017.pdf40m_REFL_RFPDs_efficiency.pdf
                      Reply  Fri May 5 10:13:02 2017, ericq, Update, General, MICH NB questions 
                         Reply  Fri May 5 12:10:53 2017, gautam, Update, General, MICH NB questions C1NB_disp_40m_MICH_NB_30_April_2017.pdf
                            Reply  Sat May 6 21:52:11 2017, rana, Update, General, MICH NB questions 
                               Reply  Wed May 10 01:56:06 2017, gautam, Update, General, MICH NB - OL coupling OL_BS_ITM_comp.pdfC1NB_disp_40m_MICH_NB_8_May_2017.pdfC1NB_disp_40m_MICH_NB_10_May_2017.pdf
                                  Reply  Wed May 10 16:53:38 2017, rana, Update, General, MICH NB - OL coupling 
                                  Reply  Wed May 10 17:17:05 2017, gautam, Update, General, DAC / Coil Driver noise actuation.jpg
                                     Reply  Wed May 10 17:46:44 2017, gautam, Update, General, DAC / Coil Driver noise - SRM coil driver + dewhite board removed 
                                        Reply  Thu May 11 09:45:46 2017, rana, Update, General, DAC / Coil Driver noise - SRM coil driver + dewhite board removed 
                                           Reply  Thu May 11 18:59:22 2017, gautam, Update, General, SRM coil driver + dewhite board initial survey D010001-B_40m.pdfD000183-C8_40m.pdf
                                              Reply  Fri May 12 01:36:04 2017, gautam, Update, General, SRM coil driver + dewhite board LISO modeling SRM_bypass_plus_CoilDriver.pdfliso.zip
                                                 Reply  Fri May 12 12:34:55 2017, gautam, Update, General, ITM and BS coil driver + dewhite board pulled out 
                                                    Reply  Fri May 12 18:50:08 2017, gautam, Update, General, ITM and BS coil driver + dewhite board pulled out 
                                                       Reply  Mon May 22 13:37:01 2017, gautam, Update, General, DAC noise estimate DAC_noise_model.pdfC1NB_disp_40m_MICH_NB_22_May_2017.pdf
                                                          Reply  Tue May 23 22:58:23 2017, gautam, Update, General, De-Whitening board noises ITMX_deWhite_ch1_noise.pdf
                                                             Reply  Thu May 25 19:27:29 2017, gautam, Update, General, Coil driver board noises  coilDriverNoises.pdf
                                                                Reply  Mon May 29 16:47:38 2017, gautam, Update, General, Coil driver boards reinstalled deWhitening_consolidated.pdfdeWhitening_measurements.zipETMY_OL.png
                                                                   Reply  Tue May 30 16:02:59 2017, gautam, Update, General, Coil driver boards reinstalled OL_noiseInjection.pdf
                                                                      Reply  Thu Jun 1 00:10:15 2017, gautam, Update, General, Coil driver boards reinstalled 
       Reply  Fri Apr 28 19:32:06 2017, gautam, Update, General, DRMI locking - PRCL angular FF PRC_angularFF.pdfPRC_TFs.pdf
Message ID: 12960     Entry time: Mon May 1 16:29:51 2017     In reply to: 12955     Reply to this: 12962
Author: gautam 
Type: Update 
Category: General 
Subject: DRMI locking  

For the traces I posted, I had not turned on the whitening for the SRCL sensing PD (REFL55). However, I took a spectrum on a subsequent lock, with the analog whitening + digital dewhitening turned on for all 3 PDs (AS55, REFL11 and REFL55), and the HF part of the SRCL spectrum still looked anomalous. I'm putting together the detailed NB, but here's a comparison between the signals from the 3 RFPDs with the PSL shutter closed (but whitening engaged, and with the analog gains at the same values as used during the locking).

 

To convert the y-axis into m/rtHz, I used data from a sensing matrix measurement I took yesterday night during a DRMI lock - I turned on lines between 300 Hz and 325 Hz for the 3DOFs for ~5 minutes, downloaded the RFPD error signal data and did the demodulation. I used numbers from this elog to convert the actuator drive from cts to m. The final numbers I used were:

MICH (AS55_Q):   8.706 * 10^11 cts/m

PRCL (REFL11_I): 2.757 * 10^12 cts/m

SRCL (REFL55_I): 1.995 * 10^10 cts/m

So it looks like there may be something weird going on with the REFL55 signal chain. Looking at the LSC rack (and also suggested by an elog search), it looks like the demodulation is done by a demod board labelled "POP55" - moreover, the demodulated outputs are taken not from the regular output ports on this board, but from the "MON" ports on the front panel. 

Quote:

one of these signals does not look like the others: explanation?

 

Attachment 1: LSC_sensingNoise.pdf  713 kB  | Hide | Hide all
LSC_sensingNoise.pdf
ELOG V3.1.3-